A solution is prepared by dissolving 25g of sucrose in 300 g of water. What is the % by mass of the solute? *




A. 6.55%



B. 9.09%



C. 7.69%



D. 8.33%

1
Expert's answer
2022-04-12T04:24:03-0400

Answer: C.


"\omega(sucrose)=\dfrac{mass\ of\ sucrose}{mass\ of\ solution}=\dfrac{25\ g}{25+300\ g}=0.0769=7.69\%"